mr.simpson drove 315 miles to get to the beach. if he averaged 45 miles per hour,how long didi the trip take?
Found 2 solutions by Alan3354, jim_thompson5910:
Answer by Alan3354(69443)

mr.simpson drove 315 miles to get to the beach. if he averaged 45 miles per hour,how long didi the trip take?
----------------
d = r*t
t = d/r
--------------
I hope he left early.

Answer by jim_thompson5910(35256)

d = rt

315 = 45t

315/45 = t

7 = t

t = 7


So it will take 7 hours.